OpenStudy (anonymous):

Can someone please explain how to factor this? d^2 + 22d + 121

OpenStudy (anonymous):

is it multiple choice

OpenStudy (anonymous):

Yes, I have the answer, but I want to know HOW to get to the answer.

OpenStudy (anonymous):

first combine like terms

OpenStudy (anonymous):

23d^3? I don't know @elitlilas

OpenStudy (anonymous):

(d+11)^2 = d^2 + 22d + 121 (d+11)(d+11) = d^2 + 22d + 121

OpenStudy (anonymous):

remember that 11*11 =121 and 11+11=22. and you cant just add exponents like that. there are no like terms here.

OpenStudy (anonymous):

I think people would benefit more from the process rather than the answer.

OpenStudy (anonymous):

I know, I wasn't asking for the answer.

OpenStudy (anonymous):

@lonnie455rich

OpenStudy (amistre64):

if it factors, then it will present itself in the form: (ax+b) (cx+d), expanding this out to compare parts ... (ax+b) (cx+d) = ac x^2 + ad x + bc x + bd = ac x^2 + (ad+bc) x + bd comparing this to the given: d^2 + 22d + 121, assume these are xs not ds x^2 + 22x + 121 ac x^2 = x^2 ; therefore ac=1 (ad+bc) x = 22x ; therefore ad + bc = 22 bd = 121 ... not much to therefore on that one if we can assume that a=c=1, this can simplify to: b+d = 22 bd = 121 if b=d=11 cant be deduced from this, then you can work it a little bit more with some substituions

OpenStudy (amistre64):

all of this really boils down to; what 2 numbers add to get the middle term (22), and multiply together to get the last term (121)

OpenStudy (anonymous):

THANK YOU! I wish someone would have told me that earlier! I've been trying to find a way to factor this expression for hours now. Giving you a medal. @amistre64

OpenStudy (amistre64):

:) youre welcome, and good luck